# MaxSup: Overcoming Representation Collapse in Label Smoothing

### v1

@@ 4.1 Intra-Class Variation and Transferability
Table 2: Metrics for feature representation quality using ResNet-50 trained on ImageNet-1K. We report intra-class variation ($\bar{d}_{\text{within}}$) and inter-class separability ($R^{2}$), both of which benefit from higher values. Although all methods reduce $\bar{d}_{\text{within}}$ relative to the baseline, MaxSup preserves the most within-class diversity.
@@
Baseline 0.3114 0.3313 0.4025 0.4451
LS 0.2632 0.2543 0.4690 0.4611
OLS 0.2707 0.2820 0.5943 0.5708
Zipf’s 0.2611 0.2932 0.5522 0.4790
MaxSup 0.2926 0.2998 0.5188 0.4972
Logit Penalty 0.2840 0.3144 0.6448 0.6024
4.2 Impact of Logit Regularization
Different regularization methods impose distinct constraints on the logit space, thereby shaping the model’s representational capacity (Kornblith et al., 2021). Among these approaches, Logit Penalty and MaxSup both act directly on logits but differ fundamentally in how they apply regularization. Logit Penalty operates by minimizing the $\ell_{2}$-norm of the entire logit vector, causing a global reduction in logit magnitudes that often induces sparsity. This uniform shrinkage can limit intra-class variation, thereby weakening the model’s ability to transfer features to downstream tasks. In contrast, MaxSup targets only the largest (top-1) logit, nudging it closer to the average logit. By selectively penalizing only the most confident prediction, MaxSup avoids universal shrinkage and preserves richer intra-class diversity, a property crucial for effective transferability. Figure 2 illustrates the distribution of logits under various regularizers. Logit Penalty yields a narrower logit range, reflecting excessive sparsity and aligning with its lower transfer performance (Table 3). By comparison, MaxSup maintains broader logit distributions, thereby retaining the fine-grained feature distinctions needed to excel on downstream tasks.
Table 3: Validation performance on CIFAR-10 with a linear probe using $l_{2}$-regularized multinomial logistic regression. Although Label Smoothing and Logit Penalty improve ImageNet accuracy, they substantially degrade transfer accuracy compared to MaxSup.
Method Linear Transfer Acc.
Baseline 0.8143
Label Smoothing 0.7458
Logit Penalty (Dauphin & Cubuk, 2021) 0.7242
MaxSup 0.8102
Figure 2: Logit density plots under three different regularization strategies: MaxSup, Logit Penalty, and standard Cross Entropy. Logit Penalty induces a narrower logit distribution, reflecting excessive shrinkage that reduces intra-class variation. By contrast, MaxSup preserves a broader range of logits and thus richer representations.
5 Experiments
@@ 5.1 Evaluation on ImageNet Classification
In this section, we assess the effectiveness of MaxSup on ImageNet-1K, comparing its performance against standard Label Smoothing and related variants.
@@ 5.1.1 Experiment Setup
Model Training Configurations
We conduct extensive experiments with both CNN and Transformer models, including the ResNet family (He et al., 2016), MobileNetV2 (Sandler et al., 2018), and DeiT-Small (Touvron et al., 2021), all thoroughly evaluated on the large-scale ImageNet dataset (Krizhevsky et al., 2012) for comprehensive performance analysis.
For ResNet Series models, we train for 200 epochs using stochastic gradient descent (SGD) with momentum 0.9, a weight decay of $1\times 10^{-4}$, and a batch size of 2048. The initial learning rate is set to 0.85 and scheduled via cosine annealing. [^2] Additional training hyperparameters follow the FFCV training scripts in https://github.com/libffcv/ffcv. See Appendix E for further details on training setups. We also evaluate ResNet-based CNNs on CIFAR-100. Here, we use an initial learning rate of 0.1, reducing it by a factor of 5 at the 60th, 120th, and 160th epochs. We train for 200 epochs with a batch size of 128, weight decay of $5\times 10^{-4}$, and Nesterov momentum set to 0.9. For DeiT-Small, we employ the official implementation and train from scratch without knowledge distillation. Although the original DeiT paper emphasizes distillation, we exclude it to provide a clearer, unbiased assessment of MaxSup’s contributions. We also omit CutMix and Mixup to retain the same optimization objective.
Hyperparameters for Compared Methods
We compare Max Suppression Regularization against multiple Label Smoothing variants, including Zipf Label Smoothing (Liang et al., 2022) and Online Label Smoothing (Zhang et al., 2021). When official implementations are available, we use them directly; otherwise, we follow the respective papers’ descriptions closely to ensure fair comparisons. All training hyperparameters are kept identical to those of the baseline models, except for algorithm-specific settings and necessary adjustments. Additionally, we employ a linearly increasing $\alpha$ scheduler, which generally benefits training and stability; see Appendix F for details. This scheduler is applied to both MaxSup and standard Label Smoothing by default to maintain consistency.
@@ 5.1.2 Experiment Results
ConvNet Comparison
Table 4 summarizes the performance of MaxSup alongside various smoothing and self-distillation methods on both ImageNet and CIFAR-100. Across all tested convolutional architectures, MaxSup achieves the highest accuracy among label smoothing–based regularizers. By contrast, OLS (Zhang et al., 2021) and Zipf-LS (Liang et al., 2022) yield less consistent gains, suggesting their reported empirical benefits may depend heavily on specific training schedules. In our reproductions of OLS and Zipf-LS, we follow the authors’ original codebases and method-specific hyperparameters but do not adopt their complete training recipes. For example, the OLS paper uses a step learning-rate scheduler over 250 epochs with an initial rate of 0.1, while Zipf-LS trains for 100 epochs under a separate set of hyperparameters. Our results underscore the robustness of MaxSup across different training setups, in contrast to the more scheme-dependent improvements noted for OLS and Zipf-LS.
@@
DeiT Comparison
Table 5 compares various regularization techniques for DeiT-Small on ImageNet. MaxSup achieves an accuracy of $76.49\%$, surpassing Label Smoothing by $0.41$ percentage points. Label Smoothing variants such as Zipf’s and OLS yield only marginally higher or comparable performance relative to standard LS, suggesting that these approaches may be less effective for vision transformer architectures—potentially due to their reliance on extensive data augmentation schemes. In contrast, MaxSup consistently outperforms both standard LS and its variants, indicating its stronger ability to enhance feature representations without additional data manipulations. These findings underscore MaxSup’s robustness across distinct model architectures, especially in settings where other regularization methods show limited effectiveness.
Table 5: Accuracy (%) comparison on DeiT-Small (Touvron et al., 2021) using different Label Smoothing variants. Results are reported as “mean ± std”; parentheses indicate absolute improvement over the baseline.
Method Accuracy (Mean) Std
@@
Table 6: Semantic segmentation results on the ADE20K validation set. Models are pretrained on ImageNet-1K and then fine-tuned with UperNet (Xiao et al., 2018) for enhanced performance. We report mean Intersection over Union (mIoU) under multi-scale (MS) testing for comprehensive evaluation.
Backbone Method mIoU (MS)
DeiT-Small (Touvron et al., 2021) Baseline 42.1
@@
Figure 3: Class activation maps generated by Grad-CAM (Selvaraju et al., 2019) for DeiT-Small models trained with MaxSup (2nd row), Label Smoothing (3rd row), and a standard Baseline (4th row). The first row shows the original images. Compared with Label Smoothing, MaxSup more effectively suppresses distractions from non-target objects and preserves key features of the target class, thereby reducing instances in which the model partially or completely focuses on irrelevant regions.
@@ 5.2 Evaluation on Semantic Segmentation
To further assess the transferability of MaxSup to downstream tasks, we evaluate its performance on semantic segmentation using the MMSegmentation framework. [^3] https://github.com/open-mmlab/mmsegmentation Specifically, we employ the UperNet architecture (Xiao et al., 2018) with a DeiT-Small backbone, trained on ADE20K. We compare backbones trained with Label Smoothing and MaxSup (on ImageNet-1K) against a baseline, following the same setup as in Section 5.1.2. During fine-tuning, all models use a standard cross-entropy loss. As shown in Table 6, MaxSup achieves a mean Intersection over Union (mIoU) of $42.8\%$, surpassing the $42.4\%$ obtained with Label Smoothing. These findings further highlight the improved feature representations afforded by MaxSup in downstream tasks such as semantic segmentation.
@@ 5.3 Visualization via Class Activation Maps
To assess how MaxSup influences model decision-making relative to Label Smoothing (LS), we employ Gradient-weighted Class Activation Mapping (Grad-CAM) (Selvaraju et al., 2019). Grad-CAM produces class-discriminative localization maps that highlight the regions most relevant to each classification decision.
We conduct comprehensive experiments on the DeiT-Small model under three distinct training setups: MaxSup (second row), Label Smoothing (third row), and standard Cross-Entropy (CE) as a baseline (fourth row). As illustrated in Figure 3, MaxSup-trained models exhibit a distinct advantage in effectively mitigating distractions caused by non-target salient objects in the background (e.g., a pole in the ‘Bird’ image, a tube in the ‘Goldfish’ image, and a cap in the ‘House Finch’ image). By contrast, LS-trained models often lose focus or erroneously attend to background objects, further reflecting the detrimental influence of LS’s Error-Enhancement term and its impact on feature learning (Please see Figures 3).
Moreover, MaxSup preserves a wider range of relevant object features, as evidenced in the ‘Shark’ and ‘Monkey’ examples, where LS-trained models fail to capture key details (fins and tails). These observations align with the analysis in Appendix G, underscoring that MaxSup better retains rich intra-class information. Consequently, MaxSup-trained models yield more accurate and reliable classifications by leveraging these detailed feature representations.
5.4 Quantitative Analysis of CAM Overlays
To further quantify the benefits of MaxSup, we measure precision and recall for Grad-CAM overlays on target object regions. Our results indicate that MaxSup consistently attains higher localization precision while maintaining or improving recall, leading to more accurate and comprehensive feature activation. Qualitative assessments similarly reveal that MaxSup-trained models exhibit both a more focused and more extensive attention map, resulting in more reliable and interpretable predictions. These findings demonstrate that MaxSup not only enhances feature representation but also contributes to more trustworthy AI systems.
@@ 6 Conclusion
In this work, we carefully examined the root causes of Label Smoothing’s (LS) shortcomings and introduced Max Suppression Regularization (MaxSup) as a targeted remedy. Our analysis shows that LS can unintentionally promote overconfidence in misclassified samples by applying insufficient regularization to erroneous top-1 logits. In contrast, MaxSup effectively addresses this by consistently penalizing the most confident logit, regardless of prediction correctness. Through extensive experiments and in-depth analyses, we demonstrate that MaxSup not only improves accuracy but also preserves richer intra-class variation and significantly enhances inter-class separability. Consequently, models trained with MaxSup capture finer-grained information about individual samples, ultimately leading to stronger transfer learning capabilities. Class activation maps further reveal that MaxSup directs model attention more accurately toward salient parts of target objects, effectively mitigating distractions from background elements.
Limitations & Future Work
Although our findings validate MaxSup’s effectiveness, several directions merit further investigation. Prior work (Müller et al., 2019) shows that teachers trained with LS can degrade performance in Knowledge Distillation (Hinton, 2015), and Guo et al. (2024) suggests LS accelerates convergence via conditioning number analysis. Future research could explore MaxSup’s impact on Knowledge Distillation workflows and its influence on training convergence. Additionally, recent studies (Súkeník et al., 2024; Garrod & Keating, 2024) indicate that $\ell_{2}$ regularization biases final layer features and weights toward lower-rank solutions than those typically associated with neural collapse. Investigating how MaxSup interacts with these low-rank biases and whether it leads to similarly optimal or novel solutions is another intriguing avenue for future work.
Impact Statement
This paper advances machine learning regularization techniques through an in-depth analysis of Label Smoothing and a proposed alternative, Max Suppression Regularization. By reinforcing generalization and transferability on widely used computer vision tasks, MaxSup has the potential to benefit a broad array of real-world applications, improving robustness without introducing known negative consequences. Thus, we anticipate a net positive impact, offering practitioners a clearer understanding of network regularization and access to more effective training methodologies.
Acknowledgements
This research was supported by the University of Washington Graduate School’s GSFEI Top Scholar Award and the Start-Up Funding awarded to Prof. Zhi-Qi Cheng. Additional support was provided by ELSA – European Lighthouse on Secure and Safe AI (No. 101070617).

### v2

@@ 4.1 Analysis of MaxSup’s Learning Benefits
Table 2: Feature quality metrics for ResNet-50 on ImageNet-1K. While all methods reduce $\bar{d}_{\text{within}}$ (intra-class variation) relative to the baseline, MaxSup retains the greatest within-class diversity. Both $\bar{d}_{\text{within}}$ and $R^{2}$ benefit from higher values.
@@
Baseline 0.311 0.331 0.403 0.445
LS 0.263 0.254 0.469 0.461
OLS 0.271 0.282 0.594 0.571
Zipf-LS 0.261 0.293 0.552 0.479
MaxSup 0.293 0.300 0.519 0.497
Logit Penalty 0.284 0.314 0.645 0.602
4.1.1 Intra-Class Variation and Transferability
As noted in Section 3.1, Label Smoothing (LS) primarily curbs overconfidence for correctly classified samples but inadvertently triggers error amplification in misclassifications. This uneven penalization can overly compress intra-class feature representations. By contrast, MaxSup uniformly penalizes the top-1 logit, whether the prediction is correct or incorrect, thereby eliminating LS’s error-amplification effect and preserving finer distinctions within each class.
Table 2 compares intra-class variation ($\bar{d}_{\text{within}}$) and inter-class separability ($R^{2}$[13] for ResNet-50 trained on ImageNet-1K. Although all investigated regularizers decrease $\bar{d}_{\text{within}}$ relative to a baseline, MaxSup yields the smallest reduction, indicating a stronger retention of subtle within-class diversity—widely associated with enhanced generalization and improved transfer performance.
These benefits are further underscored by the linear-probe transfer accuracy on CIFAR-10 (Table 3). While LS and Logit Penalty each boost ImageNet accuracy, both degrade transfer accuracy, likely by suppressing informative and transferable features. By contrast, MaxSup preserves near-baseline performance, implying that it maintains rich discriminative information crucial for downstream tasks. For extended evaluations on diverse datasets, see Table 9 in the appendix.
Table 3: Linear-probe transfer accuracy on CIFAR-10 (higher is better).
Method Acc.
Baseline 0.814
Label Smoothing 0.746
Logit Penalty 0.724
MaxSup 0.810
4.1.2 Connection to Logit Penalty
As detailed in Section 3, both Label Smoothing (LS) variants and MaxSup impose penalties directly at the logit level, aligning with the perspective that various regularizers influence a model’s representational capacity via distinct logit constraints [13]. Within this family of techniques, Logit Penalty and MaxSup both address the maximum logit, yet diverge fundamentally in their specific methods of regularization.
Logit Penalty minimizes the $\ell_{2}$-norm of the entire logit vector, inducing a global contraction that can improve class separation but also reduce intra-class diversity, potentially hindering downstream transfer. By contrast, MaxSup focuses exclusively on the top-1 logit, gently nudging it closer to the mean logit. Because only the highest-confidence prediction is penalized, MaxSup avoids the uniform shrinkage observed in Logit Penalty, preserving richer intra-class variation—a property essential for robust transfer. Further insights into this behavior can be found in Appendix N, where logit-value histograms illustrate how each method affects the logit distribution.
@@ 4.2 Evaluation on ImageNet Classification
Next, we compare MaxSup to standard Label Smoothing (LS) and various LS extensions on the large-scale ImageNet-1K dataset.
@@ 4.2.1 Experiment Setup
Model Training Configurations.We evaluate both convolutional (ResNet[11], MobileNetV2 [25]) and transformer (DeiT-Small [33]) architectures on ImageNet [15]. For the ResNet Series, we train for 200 epochs using stochastic gradient descent (SGD) with momentum0.9, weight decay of $1\times 10^{-4}$, and a batch size of 2048. The initial learning rate is 0.85 and is annealed via a cosine schedule. [^2] Additional training hyperparameters follow the FFCV scripts at https://github.com/libffcv/ffcv. See Appendix E for further details. We also test ResNet variants on CIFAR-100 with a conventional setup: an initial learning rate of 0.1 (reduced fivefold at epochs 60, 120, and 160), training for 200 epochs with batch size 128 and weight decay $5\times 10^{-4}$. For DeiT-Small, we use the official codebase [33], training from scratch without knowledge distillation to isolate MaxSup’s contribution. CutMix and Mixup are disabled to ensure the model optimization objective remains unchanged.
Hyperparameters for Compared Methods.We compare Max Suppression Regularization against a range of LS extensions, including Zipf Label Smoothing[19] and Online Label Smoothing [41]. Where official implementations exist, we adopt them directly; otherwise, we follow the methodological details provided in each respective paper. Except for any method-specific hyperparameters, all other core training settings remain identical to the baselines. Furthermore, both MaxSup and standard LS employ a linearly increasing $\alpha$-scheduler for improved training stability (see Appendix F). This ensures a fair comparison under consistent and reproducible training protocols.
@@ 4.2.2 Experiment Results
ConvNet Comparison. Table4 shows results for MaxSup alongside various label-smoothing and self-distillation methods on both ImageNet and CIFAR-100 benchmarks. Across all convolutional architectures tested, MaxSup consistently delivers the highest top-1 accuracy among label-smoothing approaches. By contrast, OLS [41] and Zipf-LS [19] exhibit less stable gains, suggesting their effectiveness may heavily hinge on specific training protocols.
To reproduce OLS and Zipf-LS, we apply the authors’ official codebases and hyperparameters but do not replicate their complete training recipes (e.g., OLS trains for 250 epochs with a step-scheduled learning rate of 0.1, and Zipf-LS uses 100 epochs with distinct hyperparameters). Even under these modified settings, MaxSup remains robust, highlighting its effectiveness across a variety of training schedules—unlike the more schedule-sensitive improvements noted for OLS and Zipf-LS.
Table 4: Performance comparison of classical convolutional networks on ImageNet and CIFAR-100. All results are shown as “mean ± std” (percentage). Bold highlights the best performance; underlined marks the second best. (Methods labeled with indicate code adapted from official repositories; see the text for additional details.)
@@
Table 5: DeiT-Small top-1 accuracy (%), reported as mean $\pm$ standard deviation. Values in parentheses indicate absolute improvements over the baseline.
Method Mean Std
@@
DeiT Comparison. Table 5 summarizes performance for DeiT-Small on ImageNet across various regularization strategies. Notably, MaxSup attains a top-1 accuracy of 76.49%, surpassing standard Label Smoothing by 0.41%. In contrast, LS variants such as Zipf-LS and OLS offer only minor gains over LS, implying that their heavy reliance on data augmentation may limit their applicability to vision transformers. By outperforming both LS and its variants without additional data manipulations, MaxSup demonstrates robust feature enhancement. These findings underscore MaxSup’s adaptability to different architectures and emphasize its utility in scenarios where conventional label-smoothing methods yield limited benefits.
Fine-Grained Classification. Beyond ImageNet, we evaluate MaxSup on demanding fine-grained benchmarks such as CUB-200 [34] and Stanford Cars [14]. As shown in Table 11, MaxSup outperforms other LS-based methods by more effectively capturing subtle within-class variations, which are critical for success in fine-grained recognition scenarios.
Ablation on the Weight Schedule. We also systematically investigate how different $\alpha$ scheduling strategies impact MaxSup’s performance. Empirical results indicate that MaxSup consistently maintains high accuracy across a wide range of schedules, further underscoring its robustness against hyperparameter changes. For additional details and discussions, refer to Appendix F.
Table 6: Semantic segmentation (multi-scale) on ADE20K using UperNet. All models are pretrained on ImageNet-1K; mIoU reported as percentage.
Backbone Method mIoU
DeiT-Small Baseline 42.1
@@ 4.3 Evaluation on Semantic Segmentation
We further investigate MaxSup’s applicability to downstream tasks by evaluating its performance on semantic segmentation using the widely adopted MMSegmentation framework. [^3] https://github.com/open-mmlab/mmsegmentation Specifically, we adopt the UperNet [36] architecture with a DeiT-Small backbone, trained on ADE20K. Models pretrained on ImageNet-1K with either MaxSup or Label Smoothing are then fine-tuned under the same cross-entropy objective (Section 4.2.2).
Table 6 shows that initializing with MaxSup-pretrained weights yields an mIoU of $42.8\%$, surpassing the $42.4\%$ achieved by Label Smoothing. This improvement indicates that MaxSup fosters more discriminative feature representations conducive to dense prediction tasks. By more effectively capturing class boundaries and within-class variability, MaxSup promotes stronger segmentation results, underscoring its potential to deliver features that are both transferable and highly robust.
Figure 2: Grad-CAM [27] visualizations for DeiT-Small models under three training setups: MaxSup (2nd row), Label Smoothing (3rd row), and a baseline (4th row). The first row shows the original images. Compared to Label Smoothing, MaxSup more effectively filters out non-target regions and highlights essential features of the target class, reducing instances where the model partially or entirely focuses on irrelevant areas.
@@ 4.4 Visualization via Class Activation Maps
To better understand how MaxSup fundamentally differs from Label Smoothing (LS) in guiding model decisions, we employ Gradient-weighted Class Activation Mapping (Grad-CAM) [27], which highlights regions most influential for each prediction.
We evaluate DeiT-Small under three training setups: MaxSup (second row), LS (third row), and a baseline with standard cross-entropy (fourth row). As illustrated in Figure 2, MaxSup-trained models more effectively suppress background distractions than LS, which often fixates on unrelated objects—such as poles in “Bird,” tubes in “Goldfish,” and caps in “House Finch.” This behavior reflects LS’s error-enhancement mechanism, which can misdirect attention.
Moreover, MaxSup retains a wider spectrum of salient features, as exemplified in the Shark” and Monkey” images, where LS-trained models often omit crucial semantic details (e.g., fins, tails, or facial contours). These findings align with our analysis in Appendix J, clearly demonstrating that MaxSup preserves richer intra-class information. Consequently, MaxSup-trained models produce more accurate and consistent predictions by effectively leveraging fine-grained object cues. Further quantitative Grad-CAM overlay metrics (e.g., precision and recall for target regions) confirm that MaxSup yields more focused and comprehensive activation maps, further underscoring its overall efficacy.
@@ 5 Conclusion
We examined the shortcomings of Label Smoothing (LS) and introduced Max Suppression Regularization (MaxSup) as a targeted and practical remedy. Our analysis shows that LS can unintentionally heighten overconfidence in misclassified samples by failing to sufficiently penalize incorrect top-1 logits. In contrast, MaxSup uniformly penalizes the highest logit, regardless of prediction correctness, thereby effectively eliminating LS’s error amplification. Extensive experiments demonstrate that MaxSup not only improves accuracy but also preserves richer intra-class variation and enforces sharper inter-class boundaries, leading to more nuanced and transferable feature representations and superior transfer performance. Moreover, class activation maps confirm that MaxSup better attends to salient object regions, reducing focus on irrelevant background elements.
Limitations. Although promising, several questions remain. Prior work[21] notes that LS-trained teachers may degrade knowledge distillation [12], and Guo et al. [9] suggests LS accelerates convergence via conditioning. Examining MaxSup’s role in distillation and its impact on training dynamics would clarify these effects. Recent studies [31, 7] also show that $\ell_{2}$ regularization biases final-layer features toward low-rank solutions, raising questions about whether MaxSup behaves similarly.
Impact. In practical applications, MaxSup shows strong promise for systems demanding robust generalization and efficient transfer, and we have not observed any additional adverse effects or trade-offs. By offering researchers and practitioners both a clearer understanding of LS’s limitations and a straightforward, computationally light, and easily integrable method to overcome them, MaxSup may help guide the development of more reliable and interpretable deep learning models.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The paper’s abstract and introduction outline the main contributions—including the identification of Label Smoothing (LS) shortcomings and the proposal of Max Suppression (MaxSup)—and these claims align with the theoretical and experimental sections.
  5. 2.
    Limitations
  6. Question: Does the paper discuss the limitations of the work performed by the authors?
  7. Answer: [Yes]
  8. Justification: A dedicated “Limitations” portion (or equivalent discussion) is provided, acknowledging possible extensions (e.g., knowledge distillation scenarios) and other open questions (e.g., interactions with $\ell_{2}$ regularization).
  9. 3.
    Theory assumptions and proofs
  10. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  11. Answer: [Yes]
  12. Justification: The paper includes formal statements and proofs in the main text and/or appendix (Lemma/Theorem with proofs in the supplementary material). All assumptions are clearly stated, and references are provided.
  13. 4.
    Experimental result reproducibility
  14. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results?
  15. Answer: [Yes]
  16. Justification: The main text and appendix provide training pipelines, hyperparameters, datasets, and references to the code. Full details (batch sizes, learning rates, etc.) are included.
  17. 5.
    Open access to data and code
  18. Question: Does the paper provide open access to the data and code?
  19. Answer: [Yes]
  20. Justification: The code is released (anonymized if needed), and the datasets used (ImageNet, CIFAR, etc.) are publicly available under their respective standard licenses.
  21. 6.
    Experimental setting/details
  22. Question: Does the paper specify all the training and test details necessary to understand the results?
  23. Answer: [Yes]
  24. Justification: Section 4.2.1 and the appendix detail the setup (optimizers, data splits, learning rates, etc.). The authors specify how they selected key hyperparameters.
  25. 7.
    Experiment statistical significance
  26. Question: Does the paper report error bars or statistical significance for the experiments?
  27. Answer: [Yes]
  28. Justification: Tables report “mean $\pm$ std” from multiple runs, reflecting the variability due to initialization or training seeds. This is shown in all main experimental tables.
  29. 8.
    Experiments compute resources
  30. Question: Does the paper provide sufficient information about compute resources?
  31. Answer: [Yes]
  32. Justification: The text or appendix indicates GPU usage (e.g., ResNet on cluster GPUs), approximate training duration, and other relevant details. Though high-level, it suffices to gauge feasibility.
  33. 9.
    Code of ethics
  34. Question: Does the research conform with the NeurIPS Code of Ethics?
  35. Answer: [Yes]
  36. Justification: The work adheres to standard academic norms, uses publicly available datasets, and presents no known ethical concerns or conflicts with the NeurIPS Code of Ethics.
  37. 10.
    Broader impacts
  38. Question: Does the paper discuss both potential positive and negative societal impacts?
  39. Answer: [Yes]
  40. Justification: The “Impact” statement addresses potential benefits (improved accuracy and transfer, leading to more robust systems) and acknowledges that misuses are minimal given the method’s purely algorithmic nature.
  41. 11.
    Safeguards
  42. Question: Does the paper describe safeguards for high-risk data or models?
  43. Answer: [N/A]
  44. Justification: The paper does not involve high-risk data (e.g., private user info) or high-risk models (e.g., generative LLMs). Standard ImageNet/CIFAR usage and training code are of no particular misuse risk.
  45. 12.
    Licenses for existing assets
  46. Question: Are the creators of assets properly credited, and the licenses mentioned?
  47. Answer: [Yes]
  48. Justification: The paper cites and credits publicly available code or datasets (ImageNet, CIFAR, etc.) with references to their original licenses or terms of service.
  49. 13.
    New assets
  50. Question: Are newly introduced assets well documented?
  51. Answer: [N/A]
  52. Justification: No new data or special code libraries are introduced beyond the regular code release. The approach modifies existing code for training frameworks but does not constitute a new dataset or model asset.
  53. 14.
    Crowdsourcing and research with human subjects
  54. Question: Are there human subjects or crowdsourcing experiments, with instructions and compensation described?
  55. Answer: [N/A]
  56. Justification: The work involves no human subjects or crowdsourcing tasks.
  57. 15.
    Institutional review board (IRB) approvals or equivalent
  58. Question: Does the paper discuss IRB approvals for human-subjects research?
  59. Answer: [N/A]
  60. Justification: The paper does not involve human subjects; no IRB is necessary.
  61. 16.
    Declaration of LLM usage
  62. Question: Does the paper describe usage of LLMs if it is essential to core methods in this research?
  63. Answer: [Yes]
  64. Justification: We used a Large Language Model (LLM) solely for writing and polishing the paper’s text. The LLM was not involved in designing, conducting, or analyzing the experiments, nor in developing the core algorithmic contributions.